OpenStudy (anonymous):

If a = -12/5, then 5x + 2y = 6 and 3x - ay = 4 are parallel.

mathslover (mathslover):

ok so first of all : \[\large{5x+2y=6}\] \[\large{5x=6-2y}\] \[\large{x=\frac{6-2y}{5}}\] right?

mathslover (mathslover):

@jordan1919 please reply

OpenStudy (anonymous):

ohh yea ok

mathslover (mathslover):

now put this value in the equation 3x-ay=4

OpenStudy (anonymous):

it doesn't need to be solved i need to figure out hot to determine if they are parallel or not.

mathslover (mathslover):

oh sorry

OpenStudy (anonymous):

no its fine.

mathslover (mathslover):

why not to find x intercept and y intercepts first?

OpenStudy (anonymous):

idk y the question is jus if it is parallel true or false i can't really understand what it's asking.

mathslover (mathslover):

ok let us try to first find x intercept what do you get?

OpenStudy (anonymous):

Firstly put a = -12/5 in the equation.. and then solve the equation..

OpenStudy (anonymous):

\[3x - \frac{-12}{5}y = 4 \implies 15x + 12y = 20\]

OpenStudy (anonymous):

You other equation is : \[5x + 2y = 6\]

OpenStudy (anonymous):

Find their slopes if they are equal then lines must be parallel.. For first one: Slope = \(\frac{-x \; coefficient}{y \; \; colefficient} = \frac{-15}{12} = \frac{-5}{4}\)

OpenStudy (anonymous):

For second one: Slope : \(\frac{-5}{2}\) As they are not equal so the line equations are not parallel..

OpenStudy (anonymous):

ohh thank for explaining like that it made it very easy to understand.